OpenStudy (anonymous):

Factor the following trinomial. w18 – 9w9y5 + 14y10

OpenStudy (anonymous):

Ummm where's the exponenets

OpenStudy (anonymous):

18, 9,5,10

OpenStudy (anonymous):

\[w ^{18}-9w ^{9}y ^{5}+14y ^{10}\] Does it look like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Calcmathlete I think you'd be better at this than me :)

OpenStudy (anonymous):

Find factors of 14 that add up to -9.

OpenStudy (anonymous):

idk that's why i'm asking for help

OpenStudy (anonymous):

so it would be -2 and -7

OpenStudy (anonymous):

Ok. Even though it looks a lot different, it is still a factorable trinomial. \[w ^{18}-9w ^{9}y ^{5}+14y ^{10} \implies (w^9 - 2y^5)(w^9 - 7y^5)\]

OpenStudy (anonymous):

Remember. THese are exponents. \[w^9 \times w^9 = w^{9 + 9} = w^{18}\]

OpenStudy (anonymous):

so would that be the answer?

OpenStudy (anonymous):

\[w ^{18}-9w ^{9}y ^{5}+14y ^{10} \implies (w^9 - 2y^5)(w^9 - 7y^5)\] The one on the right is the answer.

OpenStudy (anonymous):

thank you so much you guys!

OpenStudy (anonymous):

np :)